# Soft Condorcet Optimization for Ranking of General Agents

### v2

@@ 4.1 Warmup: Top-Ranking Condorcet Winners
We also find that Fenchel-Young gradient descent top-ranks agent $A$. This is because the gradient of a rating using the Fenchel-Young loss is weighted by the difference in ranks rather than just order misclassifications like the soft Kendall-tau distance. We elaborate on this in Section 5. Full results are shown in Appendix D.1.
4.2 Kemeny-Young Approximation Quality
@@
501 533 2190.04 50 0.46
Table 2: Kemeny-Young approximation quality on 12,680 PrefLib instances, grouped by number of alternatives where $m_{\bot}\leq|A|\leq m_{\top}$. Each row corresponds to a group, size to the number of instances in each group, $\bar{m}$ and $\bar{n}$ are the average number of alternatives and votes in each group. $C^{gd}$ and $C^{sp}$ refer to the Condorcet match proportions for gradient descent and sigmoidal programming, respectively. Similarly, $\bar{K}_{n}$ refers to normalized Kendall-tau distance to the Kemeny ranking, averaged over all instances in the group.
We evaluate approximation quality (compared to the ranking returned by the Kemeny-Young method) of Algorithm 1 on PrefLib instances (Mattei and Walsh, 2013). We run Algorithm 1 with a batch size $|B|=32$, learning rates $\alpha\in\{0.01,0.1\}$, iterations $T\in\{10^{4},10^{5}\}$, and temperature $\tau\in\{1,\frac{1}{2}$ averaged across 3 seeds per instance on all 12,680 PrefLib data instances. On the instances where $|A|\leq 10$ we also run the Kemeny-Young method. Denote an instance by $i\in\{1,2,\cdots,12680\}$. Each produces a ranking which we denote $R_{i,\text{SCO}}$ and $R_{i,\text{Kem}}$.
@@
4.3 Sparse Data Regime
In this subsection, we generate synthetic data by simulating evaluations from match-ups played in an online game setting. We do this in two ways: one that is uniform (reflecting a round-robin style tournament), and one that leads to a structured form of sparsity often encountered in competitive gaming (i.e., skill-matching platforms). We use $|A|=20$ agents where for each agent $i$: $\theta_{i}\sim N(100,30)$, and contests between 4 agents (i.e., four-player games). To generate contests, two separate distributions are used: the uniform distribution samples agents uniformly at random, and the skill-matched distribution which incrementally builds each contest, drawing 3 new candidates at random and choosing the one whose true rating is closest to the average of the set of agents so far. Then, for each contest $c$, we simulate the performance of agent $i$ in that contest, $P_{i}(c)=\theta_{i}+\epsilon_{c,i}$, where each $\epsilon_{c,i}\sim N(0,5.0)$. The outcome of the contest is then obtained by sorting the performances of all the agents in that contest. The rankings or votes are then a set of contest outcomes.
We run Algorithm 1, Elo, and several VasE methods across many simulated $n$-contest tournaments, where each value of $n$ corresponds to a proportion of missing data $p_{\emptyset}$ (alternative pairs that have not been evaluated in a contest together), with $p^{u}_{\emptyset}$ and $p^{s}_{\emptyset}$ referring the the uniform and skill-match distributions respectively (See Table 3). For each value of $n$ we run 200 instances using different seeds and report average values. For each run, we used 10000 iterations with batch size 16. We show two different metrics. First, the Kendall-tau distance between the final ranking found by Algorithm 1 and the true ranking given the true ratings (maximum value of $\frac{20\cdot 19}{2}=190$). This first metric identifies the pairs of agents whose relative order disagree between SCO and the true ranking; we denote these discordant pairs $D_{SCO,true}$. The second metric, which we call “mean true ratings distance” (MTRD), is then defined to be the average absolute difference in true ratings between all pairs of agents in these disagreements $\sum_{(i,j)\in D_{\text{SCO},\text{true}}}|\theta_{i}-\theta_{j}|/|D_{\text{SCO},\text{true}}|$, which allows us to take a nuanced look at the optimized parameters in addition to the associated ranking. The results are shown in Fig. 2.
@@ 4.3 Sparse Data Regime
Table 3: Proportion of missing data generated by uniform ($p^{u}_{\emptyset}$) and skill-matched ($p^{s}_{\emptyset}$) distributions based on the number of contests in the data set, $n$.
Figure 2: Kendall-tau distance (KTD) of ranking to true ranking, and Mean True Rating Distance (MTRD) of misranked pairs for ranking methods in tournament settings. Top: uniform distribution, Bottom: skill-matched distribution. Error bars represent 95% confidence intervals.
@@
4.4 Diplomacy Game Outcome Prediction
Figure 3: Average $\textsc{KTD}_{test}(t)$ between rankings and Diplomacy game outcomes across 50 seeds and train/test splits. All runs use batch size 32. Error bars depict 95% conf. intervals.
@@ 4.4 Diplomacy Game Outcome Prediction
We reflect the evaluation used in (Lanctot et al., 2023) with a small modification to adopt common practice in the supervised learning setting . First, we create 50 random splits of the data into training sets $\mathcal{D}_{R}$ and testing sets $\mathcal{D}_{T}$, with $|\mathcal{D}_{R}|=28049$ game outcomes (votes) and $|\mathcal{D}_{T}|=3000$ game outcomes. The random splits are such that each alternative in the test set is seen at least once on the training set, but no game outcomes (data points) are shared across train / test split. At each iteration $t$ the method has a ranking denoted $R_{t}$ learned from data in $\mathcal{D}_{R}$; we then compute and report the average Kendall-tau distance over the test set: $\textsc{KTD}_{test}(t)=\frac{1}{3000}\sum_{g\in\mathcal{D}_{T}}K_{d}(g,R_{t})$.
@@
5 Discussion
The experimental results show various properties of SCO ratings, but which algorithm should be used to compute them in general? In our experience, sigmoidal programming produced similar results to standard gradient descent using the sigmoid loss; however, it sometimes suffered from numerical instability and was hard to scale to large number of agents due to programs requiring $m^{2}$ variables.
@@ 5 Discussion
The Fenchel-Young loss is convex and hence gradient descent is guaranteed to converge to a global minimum; however, Condorcet winners (when they exist) are not necessarily top-ranked at that global minimum. Its practical performance is comparable to the sigmoid loss and slightly better in the large Diplomacy problem. Both the sigmoid and Fenchel-Young losses can be optimized online (batch size $|B|=1$) and work particularly well in approximating the optimal rankings when a large portion of evaluation data is missing and in complex ranking problem like Diplomacy.
@@
This is further illustrated in Figure 4, where we plot two landscapes of the Fenchel-Young and sigmoid losses for the same data over three agents, using the example of vote profiles in Equation 10. Since both losses are invariant by adding a constant, we plot over a 2D slice, for all values of $\theta$ with the same sum. As discussed above, the Fenchel-Young loss is strictly convex, its global minimizer exists is found by gradient descent. Like Elo, it favours the win rate and assigns the highest rating not to the Condorcet winner $C$, but to $A$. In contrast, the sigmoid loss is nonconvex, and has no global minimizer (it keeps decreasing at infinity). Optimizing over constrained rating, it assigns the highest rating to the Condorcet winner $C$.
6 Conclusion and Future Work
We introduced Soft Condorcet Optimization, a ranking framework for general agents inspired by Condorcet’s original voting system criteria; it can be applied to agent-versus-task and agent-versus-agent evaluation problems, and more generally to any problem with ordinal preference data over a set of alternatives.
The optimum of the soft Kendall-tau (sigmoid) loss is guaranteed to top-rank Condorcet winners when they exist and in practice finds one most of the time when the number of alternatives $m\leq 500$. SCO ratings are closest to the optimal ranking when data is sparse, demonstrated both in a synthetic tournament setting with ground truth ratings and also in a challenge problem with thousands of human players in the board game of Diplomacy, a setting where the running complex Condorcet methods is computationally infeasible.
For future work, we would like to compare the performance of SCO to other learning-to-rank methods, to approximations of Kemeny (Kenyon-Mathieu and Schudy, 2007; Karpinski and Schudy, 2010; Ali and Meila, 2012) or other ranking methods inspired by tournament solutions (Rajkumar et al., 2015), or alternatives (Conati et al., 2022). Also, we would like to investigate using SCO as a reward model for post-training and alignment of language models (Conitzer et al., 2024).
@@ 6 Conclusion and Future Work
Figure 4: Loss landscapes for the Fenchel-Young (left) and sigmoid (right), with minimizer (red star) over $[-1.5,1.5]^{3}$.
@@ Appendix A Proofs of Theorems
$\tilde{L}([\succeq],A,V,\bm{\theta})$ $=$ $\sum_{a,b\in A\times A}N(a,b)\sigma(\theta_{b}-\theta_{a})$
@@
$\tilde{L}([\succeq],A,V,\bm{\theta})$ $=$ $\sum_{a,b\in A\times A}N(a,b)\sigma(\theta_{b}-\theta_{a})\,.$ (11)
@@
$\tilde{L}([\succeq],A,V,\bm{\theta})$ $=$ $K+\sum_{a\succ b\in A}M(a,b)\sigma(\theta_{b}-\theta_{a}),$ (12)
@@
$\tilde{L}([\succeq],A,V,\bm{\theta})$ $=$ $K+\sum_{b\in A}M(c,b)\sigma(\theta_{b}-\theta_{c})+\sum_{a\succ b\in A_{-c}}M(a,b)\sigma(\theta_{b}-\theta_{a})\,,$ (13)
@@
Appendix B Solving a Sigmoidal Program to Compute SCO Ratings
@@
$\tilde{L}([\succeq],A,V,\bm{\theta})$ $=$ $\sum_{a,b\in A\times A}N(a,b)\sigma(\theta_{b}-\theta_{a})$
A sigmoidal program (Udell and Boyd, 2014) has the form:
@@
Appendix C Relationships between SCO and Prior Work
C.1 Relationship between SCO and Probabilistic Ranking
The problem of selecting an ordering that satisfies voters’ preferences has also been studied in statistics and psychology through probabilistic models defined over the space of permutations (Diaconis, 1988; Marden, 1995; Alvo and Yu, 2014). In particular, distance-based models (Fligner and Verducci, 1986) assume that voters have a true, collective preference over alternatives, $v^{*}$, and that the probability of observing any individual preference, $v\in V\subseteq[\succeq]$, is proportional to its distance from $v^{*}$. Thus, the probability of any individual vote $v\in V$ is
@@ C.1.1 Elo and Probablistic Rankings
A complete preference $v\in[\succeq]$ over the set of alternatives $A$ contains $\binom{|A|}{2}$ pairwise comparisons. Let the $\binom{|A|}{2}$ Bernoulli-distributed random variables $\bm{X}=\{X_{1},\ldots,X_{\binom{|A|}{2}}\}$ denote the outcome of observing every pairwise comparison on $v$. Each random variable $X_{k}$ represents the probability that alternative $a\in A$ is (strictly) preferred over alternative $b$. Thus, every Bernoulli parameter $p_{i}$ models the probability $p(a\succ b)$ as $p(x_{k}=1)=p_{k}$ and the converse preference $p(b\succ a)$ as $p(x_{k}=0)=1-p_{k}$. Pairwise comparison models (Bradley and Terry, 1952; Critchlow et al., 1991) assume that preferences between pairs of alternatives relate to their ratings $\bm{\theta}$ as:
@@
When optimized over some set of votes $v\in V\subseteq[\succeq]$ this leads to the objective:
@@
C.2 Relationship between SCO and Elo
Before we show the relation between SCO and Elo, we first summarize a well-known relationship between Elo and logistic regression (Morse, 2019).
@@ C.2 Relationship between SCO and Elo
is Elo’s predicted probability (of agent $i$ beating agent $j$), and $K=32$ (Elo, 1978). Logistic regression is a method for classification that models the prediction as a logistic function of a linear function between parameters and features (Cramer, 2002; Murphy, 2022). In binary classification, data points $(\mathbf{x},y)$ comprise input vectors $\mathbf{x}$ and classes $y\in\{0,1\}$. Define the log odds of the data point given a label of 1:
@@
How, then, does SCO relate to Elo? As in SCO, the Elo ratings are the parameters of the model and the loss is a function of the difference in ratings passed through a logistic function. The key difference is the loss function used. In Elo, the gradient update is derived from fitting a classifier to predict the probability of a win given two agents’ ratings. As such, the standard negative log-likelihood is used (Murphy, 2022). In SCO, the loss is the sum of sigmoids represented as $D_{v}$ (equation (6) and equation (5)) itself, specifically because it is interpreted as a (smooth) Kendall-tau distance– the optimal point being its minimum– rather than as a prediction.
C.3 Relationship to Perturbed Optimizers
A variety of methods have been considered to produce smooth proxys of ranking operators, i.e. functions that approximate a ranking or sorting function, that are differentiable (see, e.g. (Blondel et al., 2020b) and references therein). Indeed ranking is just one type of discrete-output operation that can be efficiently made differentiable, and we describe here the approach of (Berthet et al., 2020) following their notation: consider a general discrete optimization problem over a set $\mathcal{Y}$ that can be parameterized by an input $\bm{\theta}\in\mathbb{R}^{m}$, for a finite set of distinct points $\mathcal{Y}\subset\mathbb{R}^{m}$ with $\mathcal{C}$ being its convex hull, of the form:

### v3

@@ 4.1. Warmup: Top-Ranking Condorcet Winners
We also find that Fenchel-Young gradient descent top-ranks agent $A$. This is because the gradient of a rating using the Fenchel-Young loss is weighted by the difference in ranks rather than just order misclassifications like the soft Kendall-tau distance. We elaborate on this in Section 5. Full results are shown in Appendix E.1.
4.2. Kemeny-Young Approximation Quality
@@
501 533 2190.04 50 0.56
Table 2. Kemeny-Young approximation quality on 12,680 PrefLib instances, grouped by number of alternatives where $m_{\bot}\leq|A|\leq m_{\top}$. Each row corresponds to a group, size to the number of instances in each group, $\bar{m}$ and $\bar{n}$ are the average number of alternatives and votes in each group. $C^{gd}$ and $C^{sp}$ refer to the Condorcet match proportions for gradient descent and sigmoidal programming, respectively. Similarly, $\bar{K}_{n}$ refers to normalized Kendall-tau distance to the Kemeny ranking, averaged over all instances in the group.
We evaluate approximation quality (compared to the Kemeny-Young ranking) of Algorithm 1 on PrefLib instances Mattei and Walsh (2013). We run Algorithm 1 with a batch size $|B|=32$, learning rates $\alpha\in\{0.01,0.1\}$, iterations $T\in\{10^{4},10^{5}\}$, and temperature $\tau\in\{1,\frac{1}{2}\}$ averaged across 3 seeds per instance on all 12,680 PrefLib data instances. On the instances where $|A|\leq 10$ we also run the Kemeny-Young method. Denote an instance by $i\in\{1,2,\cdots,12680\}$. Each produces a ranking which we denote $R_{i,\text{SCO}}$ and $R_{i,\text{Kem}}$.
@@
4.3. Sparse Data Regime
In this subsection, we generate synthetic data by simulating evaluations from match-ups played in an online game setting. We do this in two ways: one that is uniform (reflecting a round-robin style tournament), and one that leads to a structured form of sparsity often encountered in competitive gaming (i.e., skill-matching platforms). We use $|A|=20$ agents where for each agent $i$: $\theta_{i}\sim N(100,30)$, and contests between 4 agents (i.e., four-player games). To generate contests, two separate distributions are used: the uniform distribution samples agents uniformly at random, and the skill-matched distribution which incrementally builds each contest, drawing 3 new candidates at random and choosing the one whose true rating is closest to the average of the set of agents so far. Then, for each contest $c$, we simulate the performance of agent $i$ in that contest, $P_{i}(c)=\theta_{i}+\epsilon_{c,i}$, where each $\epsilon_{c,i}\sim N(0,5.0)$. The outcome of the contest (vote among contestants) is then obtained by sorting the performances of all the agents in that contest.
We run Algorithm 1, Elo, and several VasE methods across many simulated $n$-contest tournaments, where each value of $n$ corresponds to a proportion of missing data $p_{\emptyset}$ (alternative pairs that have not been evaluated in a contest together), with $p^{u}_{\emptyset}$ and $p^{s}_{\emptyset}$ referring the the uniform and skill-match distributions respectively:
@@ 4.3. Sparse Data Regime
For each value of $n$ we run 200 instances using different seeds and report average values. For each run, we used 10000 iterations with batch size 16. We show two different metrics. First, the Kendall-tau distance between the final ranking found by Algorithm 1 and the true ranking given the true ratings (maximum value of $\frac{20\cdot 19}{2}=190$). This first metric identifies the pairs of agents whose relative order disagree between SCO and the true ranking; we denote these discordant pairs $D_{SCO,true}$. The second metric, which we call “mean true ratings distance” (MTRD), is then defined to be the average absolute difference in true ratings between all pairs of agents in these disagreements $\sum_{(i,j)\in D_{\text{SCO},\text{true}}}|\theta_{i}-\theta_{j}|/|D_{\text{SCO},\text{true}}|$, which allows us to take a nuanced look at the optimized parameters in addition to the associated ranking. The results are shown in Fig. 2.
Figure 2. Kendall-tau distance (KTD) of ranking to true ranking, and Mean True Rating Distance (MTRD) of misranked pairs for ranking methods in tournament settings. Top: uniform distribution, Bottom: skill-matched distribution. Error bars represent 95% confidence intervals.
@@
4.4. Diplomacy Game Outcome Prediction
@@
We reflect the evaluation used in Lanctot et al. (2023) with a small modification to adopt common practice in the supervised learning setting . First, we create 50 random splits of the data into training sets $\mathcal{D}_{R}$ and testing sets $\mathcal{D}_{T}$, with $|\mathcal{D}_{R}|=28049$ game outcomes (votes) and $|\mathcal{D}_{T}|=3000$ game outcomes. The random splits are such that each alternative in the test set is seen at least once on the training set, but no game outcomes (data points) are shared across train / test split. At each iteration $t$ the method has a ranking denoted $R_{t}$ learned from data in $\mathcal{D}_{R}$; we then compute and report the average Kendall-tau distance over the test set: $\textsc{KTD}_{test}(t)=\frac{1}{3000}\sum_{g\in\mathcal{D}_{T}}K_{d}(g,R_{t})$.
@@
Figure 3. Average $\textsc{KTD}_{test}(t)$ between rankings and Diplomacy game outcomes across 50 seeds and train/test splits. All runs use batch size 32. Error bars depict 95% conf. intervals.
5. Discussion
@@
Which algorithm should be used to compute SCO ratings? In our experience, sigmoidal programming produced similar results to standard gradient descent using the sigmoid loss; however, it sometimes suffered from numerical instability and was hard to scale to large number of agents due to programs requiring $m^{2}$ variables. Hence, we recommend using the sigmoidal programming approach only when there are a relatively low number of alternatives. The Fenchel-Young loss is convex and hence gradient descent is guaranteed to converge to a global minimum; however, Condorcet winners (when they exist) are not necessarily top-ranked at that global minimum. The practical performance of Fenchel-Young loss minimization is comparable to sigmoid loss minimization and slightly better in the large Diplomacy problem. Both the sigmoid and Fenchel-Young losses are minimized by gradient descent so can be optimized online (batch size $|B|=1$) and work particularly well in approximating the optimal rankings when a large portion of the evaluation data is missing.
@@
The differences between sigmoid and Fenchel-Young loss minimization is further illustrated in Figure 4, where we plot two landscapes of the Fenchel-Young and sigmoid losses for the same data over three agents, using the example of vote profiles in Equation 11. Since both losses are invariant by adding a constant, we plot over a 2D slice, for all values of $\theta$ with the same sum. As discussed above, the Fenchel-Young loss is strictly convex, its global minimizer exists is found by gradient descent. Like Elo, it favours the win rate and assigns the highest rating not to the Condorcet winner $C$, but to $A$. In contrast, the sigmoid loss is nonconvex, and has no global minimizer (it keeps decreasing at infinity). Optimizing over constrained rating, it assigns the highest rating to the Condorcet winner $C$.
If Condorcet-consistency or distance to the optimal ranking is important, we recommend using the sigmoid loss as it optimizes to minimize the distance to it directly; despite being non-convex, in practice it finds the Condorcet winner when it exists $\geq 96\%$ of the time when the number of alternatives $m\leq 500$ and returns almost-optimal rankings on instances where it can be compared (Section 4.2). Whereas, if assured convergence or weighting the loss functions by win rates (like Elo) is more important, we recommend using the Fenchel-Young loss optimization.
@@
Figure 4. Loss landscapes for the Fenchel-Young (left) and sigmoid (right), with minimizer (red star) over $[-1.5,1.5]^{3}$.
For future work, we would like to compare the performance of SCO to faster methods for finding Kemeny rankings (or approximations thereof) Kenyon-Mathieu and Schudy (2007); Karpinski and Schudy (2010); Ali and Meilǎ (2012); Rico et al. (2023b, a) or other ranking methods inspired by tournament solutions Rajkumar et al. (2015), or alternatives Conati et al. (2022). Considering similar differential approximations to other ranking distance functions such as Spearman’s footrule distance Diaconis and Graham (1977) as well as other ways to aggregate rankings Dwork et al. (2001) could be worthwhile. Finally, we would like to investigate using SCO and social choice theory for driving post-training for alignment of language models Conitzer et al. (2024); Swamy et al. (2024); Maura-Rivero et al. (2025).
{acks}
We would like to that thank Luke Marris and anonymous reviewers of the Social Choice and Learning Algorithms Workshop at AAMAS 2024 for their valuable feedback on earlier versions of this paper.
@@ Appendix A Proofs of Theorems
$\tilde{L}([\succ],A,V,\bm{\theta})$ $=$ $\sum_{a,b\in A\times A}N(a,b)\sigma(\theta_{b}-\theta_{a})$
@@
$\tilde{L}([\succ],A,V,\bm{\theta})$ $=$ $\sum_{a,b\in A\times A}N(a,b)\sigma(\theta_{b}-\theta_{a})\,.$ (12)
@@
$\tilde{L}([\succ],A,V,\bm{\theta})$ $=$ $K+\sum_{a\succ b\in A}M(a,b)\sigma(\theta_{b}-\theta_{a}),$ (13)
@@
$\tilde{L}([\succ],A,V,\bm{\theta})$ $=$ $K+\sum_{b\in A}M(c,b)\sigma(\theta_{b}-\theta_{c})+\sum_{a\succ b\in A_{-c}}M(a,b)\sigma(\theta_{b}-\theta_{a})\,,$ (14)
@@
Appendix B Fenchel-Young Loss Optimization
This loss, beyond having gradients that can be efficiently computed (requiring only to perform sorting on perturbed coefficients of $\theta$) is differentiable and convex in $\theta$ and can thus be efficiently over a convex set. Further, if the observed ranks $y_{i}$ are indeed generated from noisy sorting of a vector of ranks $\theta_{0}$, then the loss is minimized at $\theta_{0}$ (see Berthet et al. (2020) details).
Appendix C Solving a Sigmoidal Program to Compute SCO Ratings
@@
$\tilde{L}([\succ],A,V,\bm{\theta})$ $=$ $\sum_{a,b\in A\times A}N(a,b)\sigma(\theta_{b}-\theta_{a})$
A sigmoidal program Udell and Boyd (2014) has the form:
@@
Appendix D Relationships between SCO and Prior Work
D.1. Relationship between SCO and Probabilistic Ranking
The problem of selecting an ordering that satisfies voters’ preferences has also been studied in statistics and psychology through probabilistic models defined over the space of permutations (Diaconis, 1988; Marden, 1995; Alvo and Yu, 2014). In particular, distance-based models (Fligner and Verducci, 1986) assume that voters have a true, collective preference over alternatives, $v^{*}$, and that the probability of observing any individual preference, $v\in V\subseteq[\succ]$, is proportional to its distance from $v^{*}$. Thus, the probability of any individual vote $v\in V$ is
@@ D.1.1. Elo and Probablistic Rankings
A complete preference $v\in[\succ]$ over the set of alternatives $A$ contains $\binom{|A|}{2}$ pairwise comparisons. Let the $\binom{|A|}{2}$ Bernoulli-distributed random variables $\bm{X}=\{X_{1},\ldots,X_{\binom{|A|}{2}}\}$ denote the outcome of observing every pairwise comparison on $v$. Each random variable $X_{k}$ represents the probability that alternative $a\in A$ is (strictly) preferred over alternative $b$. Thus, every Bernoulli parameter $p_{i}$ models the probability $p(a\succ b)$ as $p(x_{k}=1)=p_{k}$ and the converse preference $p(b\succ a)$ as $p(x_{k}=0)=1-p_{k}$. Pairwise comparison models (Bradley and Terry, 1952; Critchlow et al., 1991) assume that preferences between pairs of alternatives relate to their ratings $\bm{\theta}$ as:
@@
When optimized over some set of votes $v\in V\subseteq[\succ]$ this leads to the objective:
@@
D.2. Relationship between SCO and Elo
Before we show the relation between SCO and Elo, we first summarize a well-known relationship between Elo and logistic regression Morse (2019).
@@ D.2. Relationship between SCO and Elo
is Elo’s predicted probability (of agent $i$ beating agent $j$), and $K=32$ Elo (1978). Logistic regression is a method for classification that models the prediction as a logistic function of a linear function between parameters and features Cramer (2002); Murphy (2022). In binary classification, data points $(\mathbf{x},y)$ comprise input vectors $\mathbf{x}$ and classes $y\in\{0,1\}$. Define the log odds of the data point given a label of 1:
@@
How, then, does SCO relate to Elo? As in SCO, the Elo ratings are the parameters of the model and the loss is a function of the difference in ratings passed through a logistic function. The key difference is the loss function used. In Elo, the gradient update is derived from fitting a classifier to predict the probability of a win given two agents’ ratings. As such, the standard negative log-likelihood is used Murphy (2022). In SCO, the loss is the sum of sigmoids represented as $D_{v}$ (equation (6) and equation (5)) itself, specifically because it is interpreted as a (smooth) Kendall-tau distance– the optimal point being its minimum– rather than as a prediction.
D.3. Relationship to Perturbed Optimizers
A variety of methods have been considered to produce smooth proxys of ranking operators, i.e. functions that approximate a ranking or sorting function, that are differentiable (see, e.g. Blondel et al. (2020b) and references therein). Indeed ranking is just one type of discrete-output operation that can be efficiently made differentiable, and we describe here the approach of Berthet et al. (2020) following their notation: consider a general discrete optimization problem over a set $\mathcal{Y}$ that can be parameterized by an input $\bm{\theta}\in\mathbb{R}^{m}$, for a finite set of distinct points $\mathcal{Y}\subset\mathbb{R}^{m}$ with $\mathcal{C}$ being its convex hull, of the form: